A multiple disc clutch has three discs on the driving shaft and two on the driven shaft, providing four pairs of contact surfaces. The outer diameter of the contact surfaces is 250 mm and the inner diameter is 150 mm. Determine the maximum axial intensity of pressure between the discs for transmitting 18.75 kW at 500 r.p.m. Assume uniform wear and coefficient of friction as 0.3 solution


1
Expert's answer
2021-02-25T04:15:32-0500

Let n= no. of point contact surfaces

"T=\frac{60P}{2\pi N}=\frac{60*18750}{2\pi *500}=358.09Nm"

Uniform wear

"R=\frac{r_1+r_2}{2}=\frac{125+75}{2}=100mm"

Torque

"T=n\mu w*R \implies 358090=4*0.3*w*100 \implies w=2984.08 N"

"P_{max}*r_2=c \implies c=75P_{max} N/mm"

"w=2\pi c(r_1-r_2) \implies w=2\pi *75P_{max}(r_1-r_2)
P_{max}=\frac{w}{2 \pi*75*50}"

"P_{max}=\frac{2984.08}{2 \pi*75*50}=0.127 N/mm^2"